# Training Bilingual LMs with Data Constraints in the Targeted Language

### v1

@@ 8 Limitations
Our primary focus is on English-German language training, as these two Germanic family languages share linguistic similarities Lewis et al. (2015). German is one of the most well-represented languages in the mC4 dataset, facilitating model comparisons with varying amounts of German and English data. Furthermore, the availability of extensive public resources for German, including translation systems and translated evaluation data, further supports our emphasis on this language pair. We experimented with seven additional datasets including French, Spanish, Italian, Portuguese, Korean, Japanese, and Chinese. However, we note that there are many other languages within mC4 and more broadly which can benefit from having auxiliary English data for pretraining. Due to limited evaluation benchmarks and availability of target language data for comparison, we leave investigation for truly low-resource languages to future work.
Evaluation data.
Another limitation in evaluating language models for languages other than English is that many datasets have been translated from English. These datasets may contain cultural biases or information that is not available on the web in other languages. As a result, certain aspects of the evaluation may lead to improved performance when using English auxiliary or translated data. Additionally, translated data often exhibits a distribution different from that of real data in the target languages. Therefore, an important direction for future work is the development of evaluation datasets that are not based on translation, which is essential for more accurate evaluation of multilingual language models.
Model size.
Finally, this work studies three model sizes up to 3B models. We note that there are many standard benchmarks that can be evaluated at 1B-3B scale, however many more benchmarks and patterns can appear at larger model sizes. It is important future work to evaluate whether the results extend to larger scales including evaluating potential “emergent behaviors” as well as risks at larger scales Wei et al. (2022).
@@ B.1 Train Sets
  • mC4: The primary pretraining corpus in our experiments is multilingual Colossal Clean Crawled Corpus (mC4), a curated text dataset comprising over 6.3T tokens. This corpus is derived from CommonCrawl and used for pretraining numerous language models Brown et al. (2020); Raffel et al. (2020); Touvron et al. (2023). The dataset is chosen as all languages have similar data extraction pipelines including line length filter, cld3 language detection, and deduplication Xue (2020). The English portion contains 2.7T tokens, German contains 350B tokens, French contains 320B tokens, Spanish contains 430B tokens, Portuguese contains 146B tokens, Italian contains 160B tokens, Korean contains 26B tokens, Japanese contains 160B tokens, and Chinese contains 40B tokens.
@@
  • FineWeb-EDU: A subset of the FineWeb dataset which is filtered according to a classifier trained on annotations for educational quality from Llama-3 70B model Penedo et al. (2024).
@@ B.2 Zero Shot Evaluations
Dataset EN DE FR ZH JA
ARC-C 1,172 1,137 1,147 1,146 1,147
ARC-E 2,376 2,260 2,271 2,271 2,271
HS 10,042 9,368 9,338 9,266 10,033
PIQA 1,838 1,838 1838 1,838 1,838
SCIQ 1,000 950 953 1,000 926
WG 1,267 1,184 1,215 1,059 1,096
@@ Appendix D Synthetic Prompts and Examples
Appendix E Data Ratios
We experiment with different data ratios beyond 5% used during training. For our ablation, we study the XL model (1.3B) and a 2.7B non-embedding parameters model. Results are reported in Figure 14. We see that performance is around the same for XL models, but 2.7B models perform worse with larger data ratios.
@@ Appendix E Data Ratios
Appendix F Results for Multiple Languages
F.1 Average Zero Shot Accuracy Plots
We present experimental results comparing the best performing approaches for French and German languages in Figure 10, Spanish and Portuguese in Figure 11, Italian and Korean in Figure 12, and for Japanese and Chinese languages in Figure 13. For each language, we take the best performing dataset and model found in German from Sections 3.2-5.1.
@@ F.1 Average Zero Shot Accuracy Plots
F.2 Perplexity Evaluations for Translated Training Data
@@
In Sections 5.1 and F.1 we found that performance trends were not the same across languages. In particular, French, German, Portuguese, and Spanish (belonging to the same language family) have similar patterns, however, performance for Chinese, Japanese, and Korean exhibit different patterns. To further test whether the models retain knowledge from one language in another, we translate a small portion of the training set from FineWeb-EDU and mC4 English totaling 10,000 documents. We then translate the data using the v3 translation system from Section 4.4. We measure both the macro perplexity of all documents as well as the fraction of times where the translated and original data from FineWeb-EDU (training set) have lower loss than the average loss of documents from mC4 English (not part of the training set but from a similar distribution). We refer to this quantity as translated and original exceedance. Having lower loss means the data is more familiar to the model, and having an equal exceedance across the original and translated data means the model can reason equally in either language. Our results are summarized in Table 3. We find that perplexity is nearly identical for original data, but much higher for translated data in all languages. For exceedance, in English, we see that the scores are all around 80%. However, we see that for Japanese and Chinese these values are much lower, indicating that seeing the data in English for these languages does not lower the perplexity in the target language and that the model is not making use of information in the other language. For Chinese evaluations, we note that the perplexity is much higher than for other languages indicating that the translation system potentially causes higher perplexity and lower exceedance. However, we still note that for Japanese, the exceedance is lower and expect with better translation quality, the Chinese evaluations will be similar to Japanese.
(a) 1.3B Model
(b) 2.7B Model
Figure 14: Zero-shot QA performance for 1B and 3B models at varying data ratios during training. For all models the total amount of available data in the target language is 250M tokens.
@@
F.3 Individual Eval Dataset Results
@@
Base DE $22.78\pm 1.23$ $37.92\pm 1.00$ $33.06\pm 0.47$ $62.35\pm 1.13$ $64.70\pm 1.51$ $51.14\pm 1.40$ $45.33$
Base EN $25.94\pm 1.28$ $48.15\pm 1.03$ $48.50\pm 0.50$ $71.55\pm 1.05$ $73.10\pm 1.40$ $52.01\pm 1.40$ $53.21$
Base EN + OH/ELI5 Filter $29.52\pm 1.33$ $57.03\pm 1.02$ $49.96\pm 0.50$ $71.87\pm 1.05$ $77.90\pm 1.31$ $53.83\pm 1.40$ $56.69$
ARC EN $31.83\pm 1.36$ $57.24\pm 1.02$ $48.80\pm 0.50$ $73.07\pm 1.04$ $77.00\pm 1.33$ $52.72\pm 1.40$ $56.78$
SciQ+Inst $29.10\pm 1.33$ $55.22\pm 1.02$ $48.34\pm 0.50$ $71.71\pm 1.05$ $78.50\pm 1.30$ $53.04\pm 1.40$ $55.98$
FineWeb EDU $38.14\pm 1.42$ $66.37\pm 0.97$ $54.88\pm 0.50$ $72.25\pm 1.04$ $84.60\pm 1.14$ $56.04\pm 1.39$ $62.05$
Base DE $27.44\pm 1.32$ $38.81\pm 1.03$ $39.53\pm 0.51$ $63.17\pm 1.13$ $67.68\pm 1.52$ $52.62\pm 1.45$ $48.21$
Base EN $25.07\pm 1.29$ $37.70\pm 1.02$ $36.18\pm 0.50$ $59.36\pm 1.15$ $63.05\pm 1.57$ $50.51\pm 1.45$ $45.31$
Base EN + OH/ELI5 Filter $25.77\pm 1.30$ $40.00\pm 1.03$ $35.94\pm 0.50$ $59.25\pm 1.15$ $65.05\pm 1.55$ $51.18\pm 1.45$ $46.20$
ARC EN $26.91\pm 1.32$ $39.34\pm 1.03$ $35.77\pm 0.50$ $58.65\pm 1.15$ $67.58\pm 1.52$ $52.11\pm 1.45$ $46.73$
SciQ+Inst $25.24\pm 1.29$ $39.38\pm 1.03$ $35.34\pm 0.49$ $59.74\pm 1.14$ $64.21\pm 1.56$ $52.70\pm 1.45$ $46.10$
FineWeb EDU $26.91\pm 1.32$ $42.39\pm 1.04$ $37.40\pm 0.50$ $60.45\pm 1.14$ $65.37\pm 1.54$ $50.42\pm 1.45$ $47.16$
@@
Base FR $24.66\pm 1.26$ $36.78\pm 0.99$ $32.89\pm 0.47$ $60.66\pm 1.14$ $63.90\pm 1.52$ $50.51\pm 1.41$ $44.90$
Base EN $26.28\pm 1.29$ $49.92\pm 1.03$ $48.62\pm 0.50$ $71.98\pm 1.05$ $75.20\pm 1.37$ $53.91\pm 1.40$ $54.32$
Base EN + OH/ELI5 Filter $29.52\pm 1.33$ $56.44\pm 1.02$ $49.92\pm 0.50$ $72.74\pm 1.04$ $79.70\pm 1.27$ $54.06\pm 1.40$ $57.06$
ARC EN $29.69\pm 1.34$ $57.28\pm 1.02$ $48.96\pm 0.50$ $72.91\pm 1.04$ $78.90\pm 1.29$ $53.12\pm 1.40$ $56.81$
SciQ+Inst $28.41\pm 1.32$ $53.16\pm 1.02$ $47.44\pm 0.50$ $70.51\pm 1.06$ $78.90\pm 1.29$ $54.14\pm 1.40$ $55.43$
FineWeb EDU $36.69\pm 1.41$ $65.40\pm 0.98$ $54.81\pm 0.50$ $72.74\pm 1.04$ $82.50\pm 1.20$ $55.41\pm 1.40$ $61.26$
Base FR $25.98\pm 1.30$ $38.53\pm 1.02$ $41.71\pm 0.51$ $64.25\pm 1.12$ $62.12\pm 1.57$ $53.25\pm 1.43$ $47.64$
Base EN $23.54\pm 1.25$ $37.03\pm 1.01$ $37.66\pm 0.50$ $58.98\pm 1.15$ $60.44\pm 1.58$ $49.05\pm 1.43$ $44.45$
Base EN + OH/ELI5 Filter $26.50\pm 1.30$ $38.93\pm 1.02$ $38.85\pm 0.50$ $60.50\pm 1.14$ $66.11\pm 1.53$ $50.37\pm 1.43$ $46.88$
ARC EN $25.54\pm 1.29$ $39.32\pm 1.03$ $38.17\pm 0.50$ $59.41\pm 1.15$ $64.22\pm 1.55$ $50.29\pm 1.44$ $46.16$
SciQ+Inst $25.72\pm 1.29$ $38.71\pm 1.02$ $37.61\pm 0.50$ $59.03\pm 1.15$ $63.06\pm 1.56$ $49.55\pm 1.43$ $45.61$
FineWeb EDU $27.38\pm 1.32$ $40.51\pm 1.03$ $40.67\pm 0.51$ $60.94\pm 1.14$ $64.85\pm 1.55$ $50.78\pm 1.43$ $47.52$
@@
Base ES $23.29\pm 1.24$ $38.55\pm 1.00$ $33.86\pm 0.47$ $60.83\pm 1.14$ $67.90\pm 1.48$ $51.07\pm 1.40$ $45.92$
Base EN $26.19\pm 1.28$ $48.95\pm 1.03$ $48.16\pm 0.50$ $71.11\pm 1.06$ $74.30\pm 1.38$ $52.01\pm 1.40$ $53.45$
Base EN + OH/ELI5 Filter $29.86\pm 1.34$ $58.08\pm 1.01$ $50.55\pm 0.50$ $72.36\pm 1.04$ $79.40\pm 1.28$ $54.62\pm 1.40$ $57.48$
ARC EN $30.89\pm 1.35$ $58.71\pm 1.01$ $49.41\pm 0.50$ $73.45\pm 1.03$ $78.50\pm 1.30$ $54.06\pm 1.40$ $57.50$
SciQ+Inst $29.44\pm 1.33$ $57.49\pm 1.01$ $48.89\pm 0.50$ $70.57\pm 1.06$ $79.70\pm 1.27$ $54.22\pm 1.40$ $56.72$
FWE $36.60\pm 1.41$ $64.86\pm 0.98$ $54.95\pm 0.50$ $71.87\pm 1.05$ $81.50\pm 1.23$ $57.85\pm 1.39$ $61.27$
Base ES $27.99\pm 1.33$ $44.12\pm 1.04$ $43.64\pm 0.51$ $65.94\pm 1.11$ $69.72\pm 1.49$ $52.70\pm 1.42$ $50.69$
Base EN $25.28\pm 1.28$ $38.71\pm 1.02$ $39.32\pm 0.50$ $60.94\pm 1.14$ $65.83\pm 1.54$ $47.13\pm 1.42$ $46.20$
Base EN + OH/ELI5 Filter $24.41\pm 1.27$ $35.31\pm 1.00$ $33.01\pm 0.49$ $56.64\pm 1.16$ $64.35\pm 1.55$ $50.36\pm 1.42$ $44.01$
ARC EN $25.89\pm 1.29$ $42.01\pm 1.04$ $39.49\pm 0.50$ $62.02\pm 1.13$ $66.25\pm 1.53$ $52.54\pm 1.42$ $48.03$
SciQ+Inst $25.89\pm 1.29$ $41.08\pm 1.03$ $39.35\pm 0.50$ $61.37\pm 1.14$ $67.93\pm 1.51$ $51.82\pm 1.42$ $47.91$
FWE $27.99\pm 1.33$ $43.24\pm 1.04$ $42.05\pm 0.51$ $62.19\pm 1.13$ $70.45\pm 1.48$ $51.17\pm 1.42$ $49.51$

### v2

@@ 7 Limitations
Evaluation data.
When evaluating language models for languages other than English, one often needs to rely on translated evaluation sets. Not only may this introduce translation mistakes, the resulting evaluation set might also contain cultural biases. As a result, certain aspects of the evaluation may lead to improved performance when using English auxiliary or translated data. Additionally, translated data often exhibits a distribution different from that of real data in the target languages. Therefore, an important direction for future work is the development of evaluation datasets that are not based on translation, which is essential for more accurate evaluation of bi- and multilingual language models.
@@
The focus of this work is on training language models for data constrained langauges. We note that there are many languages within mC4 (and more broadly) which can benefit from having auxiliary English data for pretraining. Due to the aforementioned limited evaluation benchmarks and availability of target language data for comparison, we leave investigation for these languages to future work.
@@ Appendix A Hyperparameters and Training Details
The 2.7B parameter model consists of 32 layers with 2560 hidden dimension and 32 attention heads. 2.7B parameter models are trained for a total of 150K steps with a batch size of 1024 and context length of 2048. All models are trained using a learning rate of $0.00016$ with the same optimization recipe as other models.
@@ B.1 Train Sets
  • mC4: The primary pretraining corpus in our experiments is multilingual Colossal Clean Crawled Corpus (mC4), a curated text dataset comprising over 6.3T tokens. This corpus is derived from CommonCrawl and used for pretraining numerous language models Brown et al. (2020); Raffel et al. (2020); Touvron et al. (2023). The dataset is chosen as all languages have similar data extraction pipelines including line length filter, cld3 language detection, and deduplication Xue (2020). The English portion contains 2.7T tokens, German contains 350B tokens, French contains 320B tokens, Spanish contains 430B tokens, Portuguese contains 146B tokens, Italian contains 160B tokens, Korean contains 26B tokens, Japanese contains 160B tokens, and Chinese contains 40B tokens. For both Chinese and Korean monolingual models: Target (L), we repeat data up to 100K steps of training (roughly 2-3 repetitions).
@@
  • FineWebEDU: A subset of the FineWeb dataset which is filtered according to a classifier trained on annotations for educational quality from Llama-3 70B model Penedo et al. (2024).
  • ChineseFineWeb-EDU [^5] https://huggingface.co/datasets/opencsg/chinese-fineweb-edu-v2: An educational corpus in Chinese consisting of roughly 400B tokens of data. Although it shares a similar name, the ChineseFineWeb-EDU does not share data from FineWebEDU and is collected from different sources. Models trained on this data are trained with only 1 repetition.
@@ B.2 Zero Shot Evaluations
We translate primarily using our own systems to (i) ensure that translation artifacts will be consistent across evaluation tasks, (ii) have translations for all languages we evaluate. We note further that evaluation frameworks such as Okapi Dac Lai et al. (2023) do not release evaluation for all languages we study, and only support HellaSwag as they do not translate the full ARC evaluation set. We did not evaluate on other multilingual evaluation tasks such as MT and MGSM as these are not traditionally knowledge tasks that require information sharing between the two languages.
Dataset EN DE FR ZH JA PT ES IT KO
ARC-C 1,172 1,137 1,147 1,146 1,147 1,117 1,117 1,117 1,147
ARC-E 2,376 2,260 2,271 2,271 2,271 2,271 2,271 2,271 2,271
HS 10,042 9,368 9,338 9,266 10,033 9,229 9,374 9,193 10,025
PIQA 1,838 1,838 1,838 1,838 1,838 1,838 1,838 1,838 1,838
SCIQ 1,000 950 953 1,000 926 953 951 952 945
WG 1,267 1,184 1,215 1,059 1,096 1,232 1,239 1,233 1,177
@@ Appendix D Synthetic Prompts and Examples
Appendix E Data Distributions
We present results comparing the data distributions for each dataset within the mC4 English clusters. The purpose of these plots is to illustrate that the data from other languages follows a similar distribution to that of mC4 English. Figure 11 shows the distribution for French and German languages, Spanish and Portuguese in Figure 12, Italian and Korean in Figure 13, and for Japanese and Chinese languages in Figure 14.
(a) German
(b) French
Figure 11: Data distribution within mC4 English clusters for German and French mC4 data.
@@ Appendix E Data Distributions
Figure 12: Data distribution within mC4 English clusters for Spanish and Portuguese mC4 data.
@@
Figure 13: Data distribution within mC4 English clusters for Italian and Korean mC4 data.
@@
Figure 14: Data distribution within mC4 English clusters for Japanese and Chinese mC4 data.
Appendix F Model Size Data Scaling
(a) 1B Model Perplexity
(b) 3B Model Perplexity
In this section, we include additional findings accompanying the results presented in §5.2.
The perplexity of training and validation data is shown in Figure 15 for the 1.3B (left) and 2.7B models (right). For 1.3B models, we see little to no overfitting at 5% ratio of training steps, and the model achieves slightly lower perplexity than a model trained with a lower ratio of 1.5% target language data. In contrast, for the 2.7B model, there is clear overfitting from as early as 25% of the training with 5% ratio of training steps. This indicates that the number of repetitions is too high for the 2.7B model and performance may degrade.
(a) 1B Model
(b) 3B Model
Figure 16: Average accuracy over zero-shot benchmark tasks in English and translated German comparing 1.3B and 2.7B models trained with 5% of the training steps using (a) 250M tokens, and (b) 1B tokens in German.
We evaluate both 1.3B and 2.7B models with 5% data ratios with 250M tokens in the target language. Results are shown in Figure 16. Figure 16(a) shows that the 2.7B model performs slightly worse than the 1.3B model with the same 250M amount of target language data matching the similar perplexity values between the two models. However, increasing to 1B tokens results in larger improvements when increasing model size from 1.3B to 2.7B, as shown in Figure 16(b).
We further conduct a comparison with different data ratios. We experiment with different data ratios beyond 5% used during training. For this experiment, we study the 1.3B parameter model and a 2.7B parameter model. Zero-shot accuracy is provided in Figure 17. We see that performance is around the same for 1.3B models, but 2.7B models perform worse with a 5% data ratio.
Appendix G Results for 300M Models
We present results comparing approaches for 300M models with English as the auxiliary language and German as the target language. Figure 17(a) shows performance with different datasets, Figure 17(b) shows performance with the DCLM filter, Figure 17(c) shows performance with upsampling real data, Figure 17(d) shows performance upsampling synthetic data, and Figure 17(e) shows performance with translation. All results show similar trends as in Section 3-4, but the performance improvements are smaller.
(a) Higher quality English auxiliary data.
(b) Model based filtering of English auxiliary data.
(c) Upsampling of English auxiliary data.
(d) Synthetic data upsampling of English auxiliary data.
(e) Translations of English auxiliary data.
Figure 17: Zero-shot accuracy of 300M models trained with different types of English auxiliary data. Results are averaged over six evaluation datasets. For each setting evaluation is done in English (pink, left bars) and German (blue, right bars).
Appendix H Results for Multiple Languages
H.1 Perplexity Evaluations for Translated Training Data
@@
In §3.3 we found that performance trends were not the same across languages. In particular, French, German, Portuguese, and Spanish (belonging to the same language family) have similar patterns, but Chinese, Japanese, and Korean exhibit different patterns. To further test whether the models retain knowledge from one language in another, we translate a small portion of the training set from FineWebEDU and mC4 English, totaling 10,000 documents. We then translate the data using the v3 translation system from §4.4. We measure both the macro perplexity of all documents as well as the fraction of times where the translated and original data from FineWebEDU (training set) have lower loss than the average loss of documents from mC4 English (not part of the training set but from a similar distribution). We refer to this quantity as translated and original exceedance. Having lower loss means the data is more familiar to the model, and having an equal exceedance across the original and translated data means the model can reason equally in either language. Our results are summarized in Table 2. We find that perplexity is nearly identical for original data, but much higher for translated data in all languages. For exceedance, in English, we see that the scores are all around 80%. However, we see that for Japanese and Chinese these values are much lower, indicating that seeing the data in English for these languages does not lower the perplexity in the target language and that the model is not making use of information in the other language. For Chinese evaluations, we note that the perplexity is much higher than for other languages, indicating that the translation system potentially causes higher perplexity and lower exceedance. However, we still note that for Japanese the exceedance is lower, and we expect that with better translation quality, the Chinese evaluations will be similar to Japanese.
@@
H.2 Other Auxiliary Languages
In this section we give additional details for our experiments with other auxiliary languages than English, as presented in §3.3.
Figure 18: Average accuracy over zero-shot benchmark tasks in translated Japanese comparing Chinese and English auxiliary data. Models are 1.3B size and trained for 100B tokens.
Methodology.
We conduct experiments comparing using English and Chinese as an auxiliary language with Japanese as the target language. For these experiments, we train a 1.3B model with the auxiliary dataset being either mC4 in English or Chinese, and English or Chinese versions of FineWebEDU [^7] https://huggingface.co/datasets/opencsg/chinese-fineweb-edu-v2. Our evaluations are on the Japanese translations.
Findings.
We summarize the results in Figure 18. Our findings indicate that using FineWebEDU, a high quality Chinese dataset, as auxiliary data improves Japanese performance over mC4 in Chinese, and the performance increases over using English either from mC4 or FineWebEDU. Note that the performance increases are consistent with improvements better English data for Indo-European langauges as seen in Figure 3.
H.3 Multilingual Experiments
Motivation.
Methodology.
We conduct experiments combining the German, French, Chinese, and Japanese language data towards multilingual training. For these experiments, we train a 1.3B model with the auxiliary dataset being FineWebEDU and a mix of data from the four languages, totaling 5% or 20% of the training. We use this subset of four languages, as we keep the data ratios the same per language, and did not want to increase the amount of data in target languages beyond the typical multilingual ratios in large open-source models Xu et al. (2024).
Findings.
EN DE FR JA ZH
5% 61.07 46.02 46.64 44.00 45.92
20% 59.25 46.44 46.61 43.23 44.36
Bi 47.16 47.52 42.73 44.38
Table 3: Evaluation of XL models in multilingual setting on “General Understanding Tasks” focusing on general reasoning, language understanding, and science knowledge in translated languages. Rows are the average accuracy for the respective language, with 5% or 20% of the training coming from a mix of the four languages. ‘Bi’ refers to the bilingual models trained with target (S) + FWE EN (L).
H.4 Average Zero Shot Accuracy Plots
We present experimental results comparing the best performing approaches for French and German languages in Figure 19, Spanish and Portuguese in Figure 20, Italian and Korean in Figure 21, and for Japanese and Chinese languages in Figure 22. For each language, we take the best performing dataset and model found in German from §3.2-3.3.
(a) German
(b) French
(a) Spanish
(b) Portuguese
(a) Italian
(b) Korean
(a) Japanese
(b) Chinese
H.5 Individual Eval Dataset Results
@@
DE (S) $20.73\pm 1.18$ $27.53\pm 0.92$ $25.94\pm 0.44$ $52.88\pm 1.16$ $36.80\pm 1.53$ $47.12\pm 1.40$ $35.17$
DE (L) $22.78\pm 1.23$ $37.92\pm 1.00$ $33.06\pm 0.47$ $62.35\pm 1.13$ $64.70\pm 1.51$ $51.14\pm 1.40$ $45.33$
DE (S) EN(L) $25.94\pm 1.28$ $48.15\pm 1.03$ $48.50\pm 0.50$ $71.55\pm 1.05$ $73.10\pm 1.40$ $52.01\pm 1.40$ $53.21$
DE (S) DCLM EN (L) $29.52\pm 1.33$ $57.03\pm 1.02$ $49.96\pm 0.50$ $71.87\pm 1.05$ $77.90\pm 1.31$ $53.83\pm 1.40$ $56.69$
DE (S) ARC EN (L) $31.83\pm 1.36$ $57.24\pm 1.02$ $48.80\pm 0.50$ $73.07\pm 1.04$ $77.00\pm 1.33$ $52.72\pm 1.40$ $56.78$
DE (S) S+I EN(L) $29.10\pm 1.33$ $55.22\pm 1.02$ $48.34\pm 0.50$ $71.71\pm 1.05$ $78.50\pm 1.30$ $53.04\pm 1.40$ $55.98$
DE (S) FWE (L) $38.14\pm 1.42$ $66.37\pm 0.97$ $54.88\pm 0.50$ $72.25\pm 1.04$ $84.60\pm 1.14$ $56.04\pm 1.39$ $62.05$
DE (S) $22.69\pm 1.24$ $30.31\pm 0.97$ $28.99\pm 0.47$ $54.84\pm 1.16$ $40.00\pm 1.59$ $50.42\pm 1.45$ $37.88$
DE (L) $27.44\pm 1.32$ $38.81\pm 1.03$ $39.53\pm 0.51$ $63.17\pm 1.13$ $67.68\pm 1.52$ $52.62\pm 1.45$ $48.21$
DE (S) EN(L) $25.07\pm 1.29$ $37.70\pm 1.02$ $36.18\pm 0.50$ $59.36\pm 1.15$ $63.05\pm 1.57$ $50.51\pm 1.45$ $45.31$
DE (S) DCLM EN (L) $25.77\pm 1.30$ $40.00\pm 1.03$ $35.94\pm 0.50$ $59.25\pm 1.15$ $65.05\pm 1.55$ $51.18\pm 1.45$ $46.20$
DE (S) ARC EN (L) $26.91\pm 1.32$ $39.34\pm 1.03$ $35.77\pm 0.50$ $58.65\pm 1.15$ $67.58\pm 1.52$ $52.11\pm 1.45$ $46.73$
DE (S) S+I EN(L) $25.24\pm 1.29$ $39.38\pm 1.03$ $35.34\pm 0.49$ $59.74\pm 1.14$ $64.21\pm 1.56$ $52.70\pm 1.45$ $46.10$
DE (S) FWE (L) $26.91\pm 1.32$ $42.39\pm 1.04$ $37.40\pm 0.50$ $60.45\pm 1.14$ $65.37\pm 1.54$ $50.42\pm 1.45$ $47.16$
@@
FR (S) $19.88\pm 1.17$ $26.60\pm 0.91$ $26.61\pm 0.44$ $52.56\pm 1.17$ $33.70\pm 1.50$ $51.14\pm 1.40$ $35.08$
FR (L) $24.66\pm 1.26$ $36.78\pm 0.99$ $32.89\pm 0.47$ $60.66\pm 1.14$ $63.90\pm 1.52$ $50.51\pm 1.41$ $44.90$
FR (S) EN(L) $26.28\pm 1.29$ $49.92\pm 1.03$ $48.62\pm 0.50$ $71.98\pm 1.05$ $75.20\pm 1.37$ $53.91\pm 1.40$ $54.32$
FR (S) DCLM EN (L) $29.52\pm 1.33$ $56.44\pm 1.02$ $49.92\pm 0.50$ $72.74\pm 1.04$ $79.70\pm 1.27$ $54.06\pm 1.40$ $57.06$
FR (S) ARC EN (L) $29.69\pm 1.34$ $57.28\pm 1.02$ $48.96\pm 0.50$ $72.91\pm 1.04$ $78.90\pm 1.29$ $53.12\pm 1.40$ $56.81$
FR (S) S+I EN(L) $28.41\pm 1.32$ $53.16\pm 1.02$ $47.44\pm 0.50$ $70.51\pm 1.06$ $78.90\pm 1.29$ $54.14\pm 1.40$ $55.43$
FR (S) FWE (L) $36.69\pm 1.41$ $65.40\pm 0.98$ $54.81\pm 0.50$ $72.74\pm 1.04$ $82.50\pm 1.20$ $55.41\pm 1.40$ $61.26$
FR (S) $22.06\pm 1.22$ $28.53\pm 0.95$ $29.25\pm 0.47$ $55.93\pm 1.16$ $40.29\pm 1.59$ $50.04\pm 1.44$ $37.68$
FR (L) $25.98\pm 1.30$ $38.53\pm 1.02$ $41.71\pm 0.51$ $64.25\pm 1.12$ $62.12\pm 1.57$ $53.25\pm 1.43$ $47.64$
FR (S) EN(L) $23.54\pm 1.25$ $37.03\pm 1.01$ $37.66\pm 0.50$ $58.98\pm 1.15$ $60.44\pm 1.58$ $49.05\pm 1.43$ $44.45$
FR (S) DCLM EN (L) $26.50\pm 1.30$ $38.93\pm 1.02$ $38.85\pm 0.50$ $60.50\pm 1.14$ $66.11\pm 1.53$ $50.37\pm 1.43$ $46.88$
FR (S) ARC EN (L) $25.54\pm 1.29$ $39.32\pm 1.03$ $38.17\pm 0.50$ $59.41\pm 1.15$ $64.22\pm 1.55$ $50.29\pm 1.44$ $46.16$
FR (S) S+I EN(L) $25.72\pm 1.29$ $38.71\pm 1.02$ $37.61\pm 0.50$ $59.03\pm 1.15$ $63.06\pm 1.56$ $49.55\pm 1.43$ $45.61$
FR (S) FWE (L) $27.38\pm 1.32$ $40.51\pm 1.03$ $40.67\pm 0.51$ $60.94\pm 1.14$ $64.85\pm 1.55$ $50.78\pm 1.43$ $47.52$
@@
ES (S) $21.16\pm 1.19$ $28.45\pm 0.93$ $26.33\pm 0.44$ $52.61\pm 1.16$ $38.50\pm 1.54$ $50.04\pm 1.41$ $36.18$
ES (L) $23.29\pm 1.24$ $38.55\pm 1.00$ $33.86\pm 0.47$ $60.83\pm 1.14$ $67.90\pm 1.48$ $51.07\pm 1.40$ $45.92$
ES (S) EN(L) $26.19\pm 1.28$ $48.95\pm 1.03$ $48.16\pm 0.50$ $71.11\pm 1.06$ $74.30\pm 1.38$ $52.01\pm 1.40$ $53.45$
ES (S) DCLM EN (L) $29.86\pm 1.34$ $58.08\pm 1.01$ $50.55\pm 0.50$ $72.36\pm 1.04$ $79.40\pm 1.28$ $54.62\pm 1.40$ $57.48$
ES (S) ARC EN (L) $30.89\pm 1.35$ $58.71\pm 1.01$ $49.41\pm 0.50$ $73.45\pm 1.03$ $78.50\pm 1.30$ $54.06\pm 1.40$ $57.50$
ES (S) S+I EN(L) $29.44\pm 1.33$ $57.49\pm 1.01$ $48.89\pm 0.50$ $70.57\pm 1.06$ $79.70\pm 1.27$ $54.22\pm 1.40$ $56.72$
ES (S) FWE (L) $36.60\pm 1.41$ $64.86\pm 0.98$ $54.95\pm 0.50$ $71.87\pm 1.05$ $81.50\pm 1.23$ $57.85\pm 1.39$ $61.27$
ES (S) $21.80\pm 1.22$ $29.90\pm 0.96$ $29.84\pm 0.47$ $57.45\pm 1.15$ $41.85\pm 1.60$ $50.61\pm 1.42$ $38.57$
ES (L) $27.99\pm 1.33$ $44.12\pm 1.04$ $43.64\pm 0.51$ $65.94\pm 1.11$ $69.72\pm 1.49$ $52.70\pm 1.42$ $50.69$
ES (S) EN(L) $25.28\pm 1.28$ $38.71\pm 1.02$ $39.32\pm 0.50$ $60.94\pm 1.14$ $65.83\pm 1.54$ $47.13\pm 1.42$ $46.20$
ES (S) DCLM EN (L) $24.41\pm 1.27$ $35.31\pm 1.00$ $33.01\pm 0.49$ $56.64\pm 1.16$ $64.35\pm 1.55$ $50.36\pm 1.42$ $44.01$
ES (S) ARC EN (L) $25.89\pm 1.29$ $42.01\pm 1.04$ $39.49\pm 0.50$ $62.02\pm 1.13$ $66.25\pm 1.53$ $52.54\pm 1.42$ $48.03$
ES (S) S+I EN(L) $25.89\pm 1.29$ $41.08\pm 1.03$ $39.35\pm 0.50$ $61.37\pm 1.14$ $67.93\pm 1.51$ $51.82\pm 1.42$ $47.91$
ES (S) FWE (L) $27.99\pm 1.33$ $43.24\pm 1.04$ $42.05\pm 0.51$ $62.19\pm 1.13$ $70.45\pm 1.48$ $51.17\pm 1.42$ $49.51$